Gio drew an angle on a coordinate grid. One side of the angle had the equation 2x - 3y = 0, and the other side had the equation 2x + 9y = 24. What were the coordinates of the vertex of the angle Gio drew?
Answer by jonvaliente(64)

To find the intersection of these two lines we need to find an x and y that is in both lines,
Subtracting 2x-3y=0 from 2x+9y=24, we get:

2x+9y=24 (1)
- 2x-3y=0 (2)
------------
12y=24
Dividing both sides by 12, we get:
y=2
Using (1) to solve for the value of x using y=2, we get:
2x+9*(2)=24
2x+18=24
2x=6
x=3
Therefore, the vertex of the angle is on the point (3,2)